Circular No. 6216

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) R. M. West, European Southern Observatory (ESO), writes: “A 30-

min red (IIIa-F + RG630) plate exposed with the ESO 1-m Schmidt telescope on Aug. 18.10 UT by G. Pizarro, J. L. Beuzit, and P. Bouchet has been photographically amplified at Garching by H.-H. Heyer. On a subtractive print (in which the difference between this and an earlier, similar plate is further enhanced), the comet is found to have a large, diffuse coma, measuring ∼ 9′.2 × 6′.0 (2.5 × 1.6 million km projected); the northeastern part is very faint. The main elongation is at p.a. ∼ 80^{o} (i.e., in the direction of the extended heliocentric radius vector). The point of maximum intensity is situated very asymmetrically, ∼ 2′ from the southwestern border of the envelope.”

D. Jewitt and J. Chen, University of Hawaii (UH), communicate: “Rband CCD images of C/1995 O1 were obtained using the UH 2.2-m telescope on Aug. 28.25 UT. A spiral arm extends from the vicinity of the nucleus at p.a. 280^{o} and curves to the north, with radial extent ∼ 8′′. Measurements of this feature may place valuable constraints on the rotation vector of the nucleus, and are encouraged.”

GRO J1735–27 S. Grebenev, K. Borozdin, and R. Sunyaev, Space Research Institute,

Moscow, on behalf of the Mir-Kvant/TTM and Granat/ART-P teams, communicate: “The ROSAT position of a new x-ray pulsar GRO J1735–27 (= RX J1735.9–2726; cf. IAUC 6207, 6212) is within the 1′ error circle for KS 1732–273, the source discovered by Mir-Kvant/TTM on 1989 Mar. 20 and Sept. 4 (in’t Zand et al. 1991, Adv. Space Res. 11, (8)187); the mean flux from the source was ∼ 22 mCrab in the band 3–10 keV during both 1989 observations. The source was marginally detected with Granat/ART-P on 1990 Apr. 4, when its flux at 3–10 keV was 13.4 ± 2.2 mCrab, but it was not found (at the flux level of 26 mCrab, 3σ upper limit) on 1990 Oct. 11. Thus, the 1989 and 1990 observations noted above, together with those on 1990 Sept. 5–7 by ROSAT and on 1995 Aug. 14 by GRO, are likely due to episodic accretion during the periastron passage (by an accreting neutron star in the binary system) and provide us with an estimate of the binary period (5–6 months).”

1995 August 29 Daniel W. E. Green